OpenStudy (anonymous):

Solve the equation x-5 over 4 = 3 for x. Please show your work and solution

OpenStudy (anonymous):

\[\frac{ x - 5 }{ 4 } = 3\]

OpenStudy (blurbendy):

(x-5) / (4) = 3 Multiply both sides by 4 x - 5 = 12 Isolate x x = 17

OpenStudy (blurbendy):

if you put 17 back in for x, (17 - 5) / 4 = 12 / 4 = 3